Identical objects are located at the same distance from two spherical mirrors, A and B. The magnifications produced by the mirrors are mA = 3.8 and mB = 1.9. Find the ratio fA/fB of the focal lengths of the mirrors.
Answer #1

Mirror equation

\frac{1}{f}=\frac{1}{d_{o}}+\frac{1}{d_{i}}

isolating the focal lenght,

f=\frac{d_{o}d_{i}}{d_{o}+d_{i}}

For mirror A,

f_{A}=\frac{d_{o}d_{iA}}{d_{o}+d_{iA}} \: \: \: \: \: (1)

For mirror B,

f_{B}=\frac{d_{o}d_{iB}}{d_{o}+d_{iB}} \: \: \: \: \: \: (2)

The magnification for both mirrors is,

M_{A}=-\frac{d_{iA}}{d_{o}}

M_{B}=-\frac{d_{iB}}{d_{o}}

or

d_{iA}=-M_{A}d_{o}

d_{iB}=-M_{B}d_{o}

inserting in eqs. (1) and (2),

f_{A}=\frac{d_{o}(-M_{A}d_{o})}{d_{o}+(-M_{A}d_{o})} =-\frac{M_{A}d_{o}}{1-M_{A}}

f_{B}=\frac{d_{o}(-M_{B}d_{o})}{d_{o}+(-M_{B}d_{o})} =-\frac{M_{B}d_{o}}{1-M_{B}}

so,

\frac{f_{A}}{f_{B}}=\frac{-\frac{M_{A}d_{o}}{1-M_{A}}}{-\frac{M_{B}d_{o}}{1-M_{B}}}=\frac{M_{A}(1-M_{B})}{M_{B}(1-M_{A})}=\frac{3.8(1-1.9)}{1.9(1-3.8)}

or

{\color{Blue} \frac{f_{A}}{f_{B}}=0,642857143=0.643}